OpenStudy (anonymous):

Which expression is equivalent to (-2)(a+6)? A. -2a+6 B. 2a+12 C. -2a-12 D. -2a+12

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

Yea!

OpenStudy (anonymous):

Alright, cool. Can you answer another few questions?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

What is the simplified form of this expression? 5(14-2)^2/2 A. 60 B. 30 C. 72 D. 360

OpenStudy (anonymous):

The answer to your second question is D. 360

OpenStudy (anonymous):

nice nice. Thank you!

OpenStudy (anonymous):

this one is an open question, so. A square field has an area of 479ft^2. What is the approximate length of a side of the field? Give your answer to the nearest foot. Explain your response.

OpenStudy (anonymous):

Area=side^2 479=side^2 \[\sqrt{479 =side}\] Does this help?

OpenStudy (anonymous):

Yes, thank you! Thanks all I needed.

OpenStudy (anonymous):

Happy I could help!